Mathematics 10 — Exercise Review · Q1 (4)
Full question
P(A)=0.6. Probability A does not occur? (a)0.4 (b)0.6 (c)1 (d)1.6
Step-by-step solution
1−0.6=0.4.
Answer
(a)0.4.
P(A)=0.6. Probability A does not occur? (a)0.4 (b)0.6 (c)1 (d)1.6
1−0.6=0.4.
(a)0.4.